• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Roedelbronn

    Log entry time 08:03:23 on March 1,2002

    Entry number 684

    keyword=Owl shift summary

    00:27 Stopped run 3374 early because beam has been down for several minutes.
    00:28 MCC called to say they were having problems with getting beam back but that it shouldn't be long.
    00:56 Beam is back and we have recommenced taking data
    01:07 Had to restart CODA because of TS2 crash, run 3375 should be good but no end-of-run log entry
    05:09 MCC taking beam for daily q.e. measurement.
    05:17 Beam is back
    08:15 Total accumulated charge for 3D ep runs thus far ~ 1.53 C